OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)2

OpenStudy (campbell_st):

well the power of 2 operates on each part of the expression which means \[(3x^3y^4)^2 = (3)^2 \times (x^3)^2 \times (y^4)^2\] the index law for power of a power is multiply the powers. \[(x^a)^b = x^{a \times b}\] use the law to simplify your expression. hope this helps

OpenStudy (anonymous):

so 27x^6y^8?

OpenStudy (campbell_st):

not quite you have to evaluate \[3^2 = \] and its not 27

OpenStudy (campbell_st):

but the rest is great work

OpenStudy (anonymous):

so its 9?